$SPAD/src/input mpoly.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool mpoly.output
)set message test on
)set message auto off
)clear all
--S 1 of 10
m : MPOLY([x,y],INT) := (x**2 - x*y**3 +3*y)**2
--R
--R
--R         4     3 3     6       2     4      2
--R   (1)  x  - 2y x  + (y  + 6y)x  - 6y x + 9y
--R                                  Type: MultivariatePolynomial([x,y],Integer)
--E 1
--S 2 of 10
m :: MPOLY([y,x],INT)
--R
--R
--R         2 6       4     3 3     2     2     4
--R   (2)  x y  - 6x y  - 2x y  + 9y  + 6x y + x
--R                                  Type: MultivariatePolynomial([y,x],Integer)
--E 2
--S 3 of 10
p : MPOLY([x,y],POLY INT)
--R
--R                                                                   Type: Void
--E 3
--S 4 of 10
p := (a**2*x - b*y**2 + 1)**2
--R
--R
--R         4 2        2   2     2      2 4       2
--R   (4)  a x  + (- 2a b y  + 2a )x + b y  - 2b y  + 1
--R                      Type: MultivariatePolynomial([x,y],Polynomial(Integer))
--E 4
--S 5 of 10
p :: POLY INT

--R
--R
--R         2 4        2          2    4 2     2
--R   (5)  b y  + (- 2a b x - 2b)y  + a x  + 2a x + 1
--R                                                    Type: Polynomial(Integer)
--E 5
--S 6 of 10
% :: MPOLY([a,b],POLY INT)
--R
--R
--R         2 4          2        2    4 2     2
--R   (6)  x a  + (- 2x y b + 2x)a  + y b  - 2y b + 1
--R                      Type: MultivariatePolynomial([a,b],Polynomial(Integer))
--E 6
--S 7 of 10
q : UP(x, FRAC MPOLY([y,z],INT))
--R
--R                                                                   Type: Void
--E 7
--S 8 of 10
q := (x**2 - x*(z+1)/y +2)**2
--R
--R
--R                             2    2
--R         4   - 2z - 2  3   4y  + z  + 2z + 1  2   - 4z - 4
--R   (8)  x  + -------- x  + ----------------- x  + -------- x + 4
--R                 y                  2                 y
--R                                   y
--RType: UnivariatePolynomial(x,Fraction(MultivariatePolynomial([y,z],Integer)))
--E 8
--S 9 of 10
q :: UP(z, FRAC MPOLY([x,y],INT))
--R
--R
--R   (9)
--R    2            3     2             2 4       3      2      2            2
--R   x   2   - 2y x  + 2x  - 4y x     y x  - 2y x  + (4y  + 1)x  - 4y x + 4y
--R   -- z  + -------------------- z + ---------------------------------------
--R    2                2                                  2
--R   y                y                                  y
--RType: UnivariatePolynomial(z,Fraction(MultivariatePolynomial([x,y],Integer)))
--E 9
--S 10 of 10
q :: MPOLY([x,z], FRAC UP(y,INT))
--R

--R
--R                                                2
--R          4      2     2  3     1  2    2     4y  + 1  2      4     4
--R   (10)  x  + (- - z - -)x  + (-- z  + -- z + -------)x  + (- - z - -)x + 4
--R                 y     y        2       2         2           y     y
--R                               y       y         y
--RType: MultivariatePolynomial([x,z],Fraction(UnivariatePolynomial(y,Integer)))
--E 10
)spool
)lisp (bye)

———-

References^{[1] nothing}

6